# Probing the limits of effective temperature consistency in actively driven systems

### v1

@@ Abstract
We investigate the thermodynamic properties of a single inertial probe driven into a nonequilibrium steady-state by random collisions with self-propelled active walkers. The probe and walkers are confined within a gravitational harmonic potential. We evaluate the robustness of the effective temperature concept in this active system by comparing values of distinct, independently motivated definitions: a generalized fluctuation-dissipation relation, a kinetic temperature, and a work fluctuation relation. Our experiments reveal that, under specific conditions, these independent measurements yield a remarkably consistent effective temperature over a wide range of system configurations. Furthermore, we also identify regimes where this consistency breaks down, which delineates the fundamental limits of extending equilibrium-like thermodynamic concepts to athermal, actively driven systems.
@@ I Introduction
Near and at equilibrium, $T$ can be equivalently defined through various thermodynamic relations, which can be applied to and measured by a colloidal thermometer. Static relations, measured at stationary states, include canonical distributions and energy equipartition, as well as configurational definitions of temperature [4, 5]. Dynamic relations, measured by probing the system’s response properties, are given by the fluctuation-dissipation theorem [6, 7, 8, 9, 10], which rigorously connects the effects of thermal noise and energy dissipation. In addition, fluctuation theorems also relate temperature to fluctuations in thermodynamic quantities such as heat, work, and entropy [11, 12, 13].
@@
Colloidal thermometers have thus been widely employed in experiments, both to validate theoretical measures of equilibrium temperature and to identify deviations from it [14, 15, 16, 17]. Crucially, various definitions of absolute temperature must yield consistent, measurement-independent values in accord with the zeroth law of thermodynamics.
@@
In stark contrast, no universal definition for temperature exists for many natural and engineered systems that operate far from thermal equilibrium. Nonetheless, the notion of effective temperature has frequently arisen as a heuristic extension of equilibrium statistical mechanics to diverse systems operating far from equilibrium, including glassy materials [18, 19, 20, 21], biological suspensions [22, 14, 15, 23, 24, 25, 26], driven granular media [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], and active matter [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 17, 56].
@@
In particular, under well-separated timescales, an equilibriumlike fluctuation-dissipation relation (FDR) can become applicable, leading to a definition of an effective temperature, $T_{\text{eff}}$. While the FDR-based $T_{\text{eff}}$ has proven useful in various fields of physics[57, 58, 16, 59, 60], its thermodynamic interpretation and range of applicability remain subjects of ongoing research [61, 62, 63, 64].
A remaining key challenge lies in identifying under which conditions distinct theoretical definitions of effective temperatures yield mutually consistent values far from equilibrium [65, 66]. Although numerous thermodynamic-like criteria have been proposed to establish a meaningful $T_{\text{eff}}$ for nonequilibrium steady-states (NESS), consistency across measurements remains largely supported by empirical evidence [15, 40].
Figure 1: The system:  A. Experimental setup: A Styrofoam ball (diameter $\sim 4$ cm, 11 gr) is trapped in a gravitational harmonic potential, a plastic bowl (diameter 38cm, depth 5cm), and subjected to collisions with $N_{b}=10$ self-propelled bbots (inset: standard bbot, $4\times 1$ cm, $7.1$ gr). The ball is repeatedly perturbed with a uniform air stream created by an external fan along the $x$-axis (white arrow) to test a fluctuation-response relation. To enforce an abrupt onset and release of the perturbation, a mechanical shutter is used (denoted by ’S’).  B. Velocity distributions of bbot assemblies with different $N_{b}=3,4,$ and 10. Instantaneous velocities were extracted from 30 fps recordings of bbot trajectories tracked over 40 minutes.  C. Spatial distribution of 10 bbots within the harmonic trap, with a soft boundary (steep curvature) indicated by a dashed circle of radius $10$ cm.  D, E. Typical 2020 second trajectories of a single bbot in systems with $N_{b}=3$ (D) and $N_{b}=10$ (E) (upper panels), alongside the corresponding tracer trajectories (lower panels).
@@
In this work, a statistical thermometer is realized experimentally, based on a macroscopic tracer particle confined in a harmonic trap and driven into a NESS by random inelastic collisions with self-propelled walkers (see Section 2 and Fig. 1). Previously [67], it was shown that this system satisfies a linear (dynamic) FDR, with a large enough number of walkers. Here, we assess the consistency of the FDR-based $T_{\text{eff}}$ by comparing it to three independent measures: the tracer’s potential energy and modified kinetic energy (at an unperturbed NESS), and a temperature based on a steady-state work fluctuation relation (FR), in a strongly perturbed NESS. Our main result is displayed in Fig. 2: all three independent measurements of $T_{\text{eff}}$ agree across a range of active bath parameters. We further compare our results with theoretical predictions and discuss conditions under which the consistency of $T_{\text{eff}}$ breaks down.
@@ II Experimental setup and basic characterization
An external airflow can be applied (in the $x$ direction) and abruptly switched-off by a physical shutter, affecting primarily the tracer particle. This mechanism is used to perturb the tracer’s NESS and measure a response function, as detailed in Section 3.1.
@@
Nonequilibrium properties of the active bath.-  In this setting, a single bbot with sufficient inertia typically performs circular motion around the trap center [68], with a preferred chirality (clockwise for Hexbugs). As the number of bbots $N_{b}$ increases, frequent collisions with other bbots (and the tracer) randomize their propulsion direction and result in an active gas-like state (see SM movie 1). In particular, the bbot velocity distribution remains independent of $N_{b}$ whereas its position distribution deviates from the original circular path with large $N_{b}$, as seen in Fig. 1B,C.
@@
Typical bbot trajectories are shown in Fig. 1D,E. For $N_{b}=3$, collisions are rare, and bbots follow long-lived circular paths; for $N_{b}=10$, frequent collisions yield erratic, stochastic motion. Notably, these self-propelled particles exhibit rich dynamics, including alignment with boundaries and forces and emergent collective motion [69]. In contrast, the passive tracer primarily experiences random inelastic collisions with the active bbots (lower panels). These interactions introduce both noise and dissipation, which, in the spirit of the Langevin framework, result in a Brownian-like motion within the harmonic potential. In this work, we focus on the properties of the tracer particle, as detailed below.
@@
Nonequilibrium properties of the tracer.-  The NESS statistics and dynamics of the tracer particle (in unperturbed conditions) are shown in Fig. 3, with the number of bbots $N_{b}$ used as a control parameter. These results were obtained by averaging 375 particle trajectories of 1-minute duration and a time interval of $\Delta t=1/30$ s.
Both the stationary position ($x$) and velocity ($v_{x}$) distributions of the tracer generally deviate from Gaussian behavior. As shown in Fig. 3A, the position distribution evolves with increasing $N_{b}$, transitioning from an exponential form toward a Gaussian-like profile, particularly within the harmonic trap boundary $|x|<10$ cm (see also Ref. [67]). The velocity distributions (Fig. 3B) exhibit an exponential decay for small $N_{b}=3$, with reliable statistics within $|v_{x}|<50$ cm/s. At larger $N_{b}$, the distributions converge to a distinct non-Gaussian shape that is well described by a stretched exponential function.
The stationary dynamics of the tracer are characterized in Figs. 3C,D via the position and velocity autocorrelation functions (ACF). An increase in the rate of dissipative collisions upon adding more bbots, leads to more strongly damped tracer motion. This is reflected in a suppression of oscillations in the relaxation dynamics, and in the reduction of the relaxation time.
@@
Figure 2: Thermometer consistency:  $N_{b}=\{3,4,6,10,12,15,17\}$ bbots. A. Effective temperature measurements obtained using three independent methods: the potential temperature $T_{\text{pot}}=T_{\text{eff}}$ (circles, Eq. 4), the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (squares, Eq. 6), and a constant temperature $T_{\text{FR}}$ derived from a work FR (diamonds, Eq. 7). Notably, these static temperatures validate the FDR of Eq. 2 (for $N_{b}>3$) and define a consistent effective temperature $T_{\text{eff}}$.  B. Exemplary results for three independent steady-states, used for effective temperature measurements under different conditions: (a) an unperturbed state yielding stationary kinetic and potential temperatures ($T_{\text{pot}}$ and $\tilde{T}_{\text{kin}}$); (b) weakly perturbed state characterized by a linear fluctuation-response relation and its associated $T_{\text{eff}}$; and (c) a strongly perturbed state with a biased work distribution yielding $T_{\text{FR}}$ via a fluctuation relation.
@@
where the bar indicates normalization, $\langle v_{x}^{2}\rangle_{0}$ is the (stationary) velocity variance, and $\omega^{2}\equiv\Omega^{2}-\gamma^{2}/4$. Fitting the ACF to Eq. 1 provides system-specific values of the effective timescales: the damping rate, $\gamma$, and the harmonic frequency, $\Omega$.
These timescales are displayed in Fig. 3F as a function of $N_{b}$, showing a transition between underdamped ($\Omega<\gamma/2$) towards critically-damped ($\Omega\approx\gamma/2$) dynamics as $N_{b}$ increases. A reduction in the relaxation time, $2\gamma^{-1}$, corresponds to an increase in collision frequency, evaluated by image analysis as the mean-free time between tracer-bbot collisions $\tau_{c}$ (black line).
At high densities ($N_{b}=15$ and $17$), fast sequential collisions require a higher temporal resolution to capture the tracer’s inertial dynamics. In these cases, the fit parameters are consistent with the critical regime under $\Delta t=1/60$ s, whereas $\Delta t=1/30$ s results in overdamped behavior. These measurements of velocity fluctuations and their corresponding system-specific timescales are further discussed in the context of the kinetic temperature in Sec. 3.2.
@@ III Effective temperatures
Figure 3: NESS dynamics and statistics: Data were obtained for active bath configurations with $N_{b}=\{3,4,6,10,12,15,17\}$ bbots, using a tracer of mass $m=1\pm 0.1$ g confined in a potential of stiffness $\kappa=28.2\pm 3$ g/s2. Results present averages over 375 one-minute tracer trajectories recorded at 30 fps for each $N_{b}$.  A. Normalized position ($x$) distributions. Dashed line show fits to exponential ($N_{b}=3$) and Gaussian ($N_{b}=15$) functions.  B. Normalized velocity ($v_{x}$) distributions. An exponential fit is shown for $N_{b}=3$ and a stretched exponential fit for $N_{b}=15$.  C,D. Position (C) and velocity (D) autocorrelation functions (ACFs), showing increasingly damped dynamics with higher $N_{b}$.  E. Position and velocity ACFs for $N_{b}=10$, fitted with Eq. 1 using $\gamma$ and $\Omega$ as fitting parameters, yielding consistent descriptions for both $C_{xx}$ and $C_{vv}$.  F. Extracted relaxation rate ($\gamma$) and trapping frequency ($\Omega$) as functions of $N_{b}$ (upper panel). The evaluated mean-free time between tracer-bbot collisions ($\tau_{c}$) is plotted as frequency versus $N_{b}$ (lower panel).
@@ III.1 FDR temperature
where $R_{x}$ and $R_{v}$ are the mean response functions of the position and velocity observables, respectively. The correlation functions in the (long-time) unperturbed NESS are given by a position ACF, $C_{xx}=\langle x(t)x(0)\rangle_{0}$, and a position-velocity cross-correlation function, $C_{vx}=\langle v(t)x(0)\rangle_{0}$. In this context, an effective temperature $T_{\text{eff}}$ is defined as a proportionality constant.
Notably, a static FDR can be recovered in Eq. 2 at $t=0$, with $C_{xx}(0)\rightarrow\langle x^{2}\rangle_{0}$ and $R_{x}(0)\rightarrow\Delta x_{\epsilon}$, where $\langle x^{2}\rangle_{0}$ is the position variance in the unperturbed NESS, and $\Delta x_{\epsilon}$ in the mean displacement. Considering that $F_{0}=\kappa\Delta x_{\epsilon}$, the effective temperature $T_{\text{eff}}$ coincides with the definition of potential energy equipartition,
@@
Thus, implying that $T_{\text{eff}}=T_{\text{pot}}$ can be defined if the dynamic FDR in Eq. 2 is validated, even in the presence of non-Gaussian steady-state distributions.
@@
Fig. 4C presents the corresponding FDR tests for both position and velocity observables. The system’s full response to the abrupt force arrest at time $t_{1}=t-t_{0}$ is evaluated as $R_{x}(t_{1})=\langle x(t_{1})\rangle-\langle x\rangle_{0}$ and $R_{v}(t_{1})=\langle v(t_{1})\rangle-\langle v\rangle_{0}$, and compared with the unperturbed correlation functions, $C_{xx}$ and $C_{xv}$. Within measurement error, these results validate the FDRs given in Eqs. 2 and 3, with an effective temperature determined by Eq. 4.
Thereby, in this system the tracer’s potential temperature satisfies a dynamic FDR in both position and velocity observables. We note that the signal-to-noise ratio in $R_{v}$ diminishes with increasing $N_{b}$, rendering the velocity response experimentally inaccessible for larger values ($N_{b}>10$). The independent FDR tests of the experimental setups are further provided in the SI.
Fig. 4D presents a parametric plot of the position FDR, showing normalized fluctuation and response quantities across different $N_{b}$ configurations. Specifically, the solid linear line represents validation of Eq. 2 with $T_{\text{eff}}=T_{\text{pot}}$. Violations are mainly observed with the lowest $N_{b}=2$ and $3$, indicating a breakdown of the effective temperature description (as detailed in Ref. [67]). In contrast, with larger $N_{b}$, the tracer’s potential temperature $T_{\text{pot}}$ satisfies a dynamic FDR and is therefore consistent with $T_{\text{eff}}$.
Figure 4: FDR test of the full-response to a step-perturbation:  Results present averages over 375 two-minute step-perturbation sequences (at $t_{0}=60$ s), under weak external airflow applied by a fan at an operating voltage 99 V. A. Example of a typical perturbation sequence $x(t)$, where the fan is turned on for one minute and abruptly turned off for the subsequent minute.  B. Time-dependent mean position $\langle x(t)\rangle$ and mean velocity $\langle v(t)\rangle$, obtained within a two-minute time window.  C. FDR analysis using Eq. 2 (lower panel) and Eq. 3 (upper panel). These results were obtained with $N_{b}=10$.  D. Parametric plot of position FDRs for $N_{b}=\{2,3,4,6,10,12,15,17\}$ realizations, where the effective temperature $T_{\text{eff}}$ determines the linear slope. In equilibrium-like behavior, all FDR data collapse onto a single line with slope $T_{\text{eff}}\sim\kappa\langle x^{2}\rangle_{0}$. Clear deviations from this linearity are observed for the lowest densities ($N_{b}=2$ and 3, triangles), indicating FDR violations.

### v2

@@ Abstract
We investigate the thermodynamic properties of a single inertial probe driven into a nonequilibrium steady state by random collisions with self-propelled active walkers. The probe and walkers are confined within a gravitational harmonic potential. We evaluate the robustness of the effective temperature concept in this active system by comparing values of distinct, independently motivated definitions: a generalized fluctuation-dissipation relation, a kinetic temperature, and via a work fluctuation relation. Our experiments reveal that, under specific conditions, these independent measurements coincide over a wide range of system configurations, yielding a remarkably consistent effective temperature. Furthermore, we also identify regimes where this consistency breaks down, which delineates the fundamental limits of extending equilibrium-like thermodynamic concepts to athermal, actively driven systems.
@@ I Introduction
Near and at equilibrium, $T$ can be equivalently defined through various thermodynamic relations, which can be applied to and measured by a colloidal thermometer. Static relations, measured in stationary states, include canonical distributions and energy equipartition, as well as configurational definitions of temperature [39, 70]. Dynamic relations, measured by probing the system’s response properties, are given by the fluctuation-dissipation theorem [23, 48, 11, 46, 47], which rigorously connects the effects of thermal noise and energy dissipation. In addition, fluctuation theorems also relate temperature to fluctuations in thermodynamic quantities such as heat, work, and entropy [29, 28, 42, 36].
@@
Colloidal thermometers have thus been widely employed in experiments, both to validate theoretical measures of equilibrium temperature and to identify deviations from it [49, 54, 16, 35]. Crucially, various definitions of absolute temperature must yield consistent, measurement-independent values in accordance with the zeroth law of thermodynamics.
@@
In stark contrast, no universal definition for temperature exists for many natural and engineered systems that operate far from thermal equilibrium. Nonetheless, the notion of effective temperature has frequently arisen as a heuristic extension of equilibrium statistical mechanics to diverse systems operating far from equilibrium, including glassy materials [17, 60, 5, 43, 8], biological suspensions [86, 49, 54, 2, 53, 84, 20], driven granular media [66, 64, 79, 59, 81, 80, 74, 10, 65, 25, 1, 44, 34, 15, 89], and active matter [61, 57, 50, 21, 31, 63, 26, 27, 38, 75, 88, 62, 9, 72, 35, 30].
@@
In particular, under well-separated timescales, an equilibriumlike fluctuation-dissipation relation (FDR) can become applicable, leading to a definition of an effective temperature, $T_{\text{eff}}$. While the FDR-based $T_{\text{eff}}$ has proven useful in various fields of physics[55, 3, 16, 52, 4], its thermodynamic interpretation and range of applicability remain subjects of ongoing research [56, 18, 67, 41, 76].
[…]
A remaining key challenge lies in identifying under which conditions distinct theoretical definitions of effective temperatures yield mutually consistent values far from equilibrium [14, 40]. Further examples of nonequilibrium definitions that capture meaningful physical behavior in their statistical formulation include equipartition-based approaches, such as granular temperatures [64, 66], and those derived from fluctuation-relations (FR), applicable to both long-time nonequilibrium steady states (NESS) and nonequilibrium transitions [28, 25, 68]. Such definitions typically require stable steady states, the coexistence of stochastic and deterministic dynamics, and access to distinct NESS observables - conditions that are often difficult to achieve experimentally. Although no theoretical principle ensures the coincidence of these definitions far from equilibrium, our experiments indicate a range of conditions in which a consistent effective temperature can emerge in NESS, driven by active or athermal fluctuations (cf. [60, 54, 15]).
@@
In this work, a statistical thermometer is realized experimentally, based on a macroscopic tracer particle confined in a harmonic trap and driven into a NESS by random inelastic collisions with self-propelled walkers (see Section 2 and Fig. 1). Previously [7], it was shown that, with a large enough number of walkers, this system satisfies a linear (dynamic) FDR. Here, we assess the consistency of the FDR-based $T_{\text{eff}}$ by comparing it to three independent measures: the tracer’s potential energy and modified kinetic energy (in an unperturbed NESS), and a temperature based on a steady state work FR, in a NESS under a strong external perturbation. Our main result is displayed in Fig. 2: all three independent measurements of $T_{\text{eff}}$ agree across a range of active bath parameters. We further compare our results with theoretical predictions and discuss conditions under which the consistency of $T_{\text{eff}}$ breaks down.
@@ II Experimental setup and basic characterization
An external airflow can be applied via a fan (in the $x$ direction) and abruptly switched-off by a physical shutter, affecting primarily the tracer particle. Stationary probability densities of the tracer’s position along the $x$-axis for different intensities of air streams are shown in Fig. 1B. The main effect of the air stream is a shift of the mean value of the histograms in the potential well, i.e., an unperturbed state (a), a weakly perturbed state (b), and a strongly perturbed state (c). This mechanism is used to test the validity of both FDRs and FRs, as detailed in Sections 3.1 and 3.3 respectively (see SI figures 1-2 for further details). We assume independent statistics along all axes and focus on analyzing the tracer’s position and velocity components projected onto the $x$-axis, where the perturbation occurs.
@@
Nonequilibrium properties of the active bath.- In this setting, a single bbot with sufficient inertia typically performs circular motion around the trap center [19], with a preferred chirality (clockwise for Hexbugs). As the number of bbots $N_{b}$ increases, frequent collisions with other bbots (and the tracer) randomize their propulsion direction and result in an active gas-like state (see SM movie 1). In particular, the bbot velocity distribution remains independent of $N_{b}$ whereas its position distribution deviates from the original circular path with large $N_{b}$, as seen in Fig. 1C,D.
@@
Typical bbot trajectories are shown in Fig. 1E,F. For $N_{b}=3$, collisions are rare, and bbots follow long-lived circular paths; for $N_{b}=10$, frequent collisions yield erratic, stochastic motion. In general, these self-propelled particles exhibit rich dynamics, including alignment with boundaries and emergent collective motion [32]. Here, the passive tracer primarily experiences random inelastic collisions with the active bbots (lower panels). These interactions introduce both noise and dissipation, which result in a collision-induced stochastic motion reaching a steady state within the harmonic potential. In this work, we focus on the properties of the tracer particle, as detailed below.
@@
Nonequilibrium properties of the tracer.- The NESS statistics and the dynamics of the tracer particle (in unperturbed conditions) are shown in Fig. 3, with the number of bbots $N_{b}$ used as a control parameter. These results were obtained by averaging 375 particle trajectories of 1-minute duration and a time interval of $\Delta t=1/30$ s.
@@
Figure 2: Thermometer consistency: $N_{b}=\{3,4,6,10,12,15,17\}$ bbots. Effective temperature measurements obtained using three independent methods: the potential temperature $T_{\text{pot}}=T_{\text{eff}}$ (circles, Eq. 4), the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (squares, Eq. 6), and a constant temperature $T_{\text{FR}}$ derived from a work FR (diamonds, Eq. 7). Notably, these static temperatures validate the FDR of Eq. 2 (for $N_{b}>3$) and define a consistent effective temperature $T_{\text{eff}}$.
Figure 3: NESS dynamics and statistics: Data were obtained for active bath configurations with $N_{b}=\{3,4,6,10,12,15,17\}$ bbots, using a tracer of mass $m=1\pm 0.1$ g confined in a potential of stiffness $\kappa=28.2\pm 3$ g/s2. Results are obtained from combined time and ensemble averages over an ensemble of 375 one-minute-long tracer trajectories recorded at 30 fps for each $N_{b}$. A. Rescaled position ($x$) distributions. Dashed line show fits to exponential ($N_{b}=3$) and Gaussian ($N_{b}=15$) functions. B. Rescaled velocity ($v_{x}$) distributions. An exponential fit is shown for $N_{b}=3$ and a stretched exponential fit for $N_{b}=15$. C,D. Position (C) and velocity (D) autocorrelation functions (ACFs), showing increasingly damped dynamics with higher $N_{b}$. E. Position and velocity ACFs for $N_{b}=10$, fitted with Eq. 1 using $\gamma$ and $\Omega$ as fitting parameters, yielding consistent descriptions for both $C_{xx}$ and $C_{vv}$. F. Extracted relaxation rate ($\gamma$) and trapping frequency ($\Omega$) as functions of $N_{b}$ (upper panel). The impact of under-sampling on the observed dynamics is pronounced for large $N_{b}$ ($15$ and $17$). Hollow markers (dashed lines) show measurements at $30$ fps, while colored markers correspond to $60$ fps. The lower frame rate leads to apparent overdamped behavior, whereas the higher frame rate reveals dynamics consistent with a critically-damped regime. The collision frequency ($\tau_{c}^{-1}$) is plotted versus $N_{b}$ (lower panel), where $\tau_{c}$ is the measured mean-free time between tracer-bbot collisions.
Both the stationary position ($x$) and velocity ($v_{x}$) distributions of the tracer generally deviate from Gaussian behavior. As shown in Fig. 3A, the position distribution evolves with increasing $N_{b}$, transitioning from an exponential form toward a Gaussian-like profile, particularly within the harmonic trap boundary $|x|<10$ cm (for $|x|>10$ cm the arena possesses a steeper curvature, i.e., soft boundary for the trapped tracer and bbots). The velocity distributions (Fig. 3B) exhibit an exponential decay for small $N_{b}=3$, with reliable statistics within $|v_{x}|<50$ cm/s. At larger $N_{b}$, the velocity distributions converge to a distinct non-Gaussian shape that is well described by a stretched exponential function.
The stationary dynamics of the tracer are further characterized in Figs. 3C,D via the position and velocity autocorrelation functions (ACF). An increase in the rate of dissipative collisions upon adding more bbots, leads to more strongly damped tracer motion. This is reflected in a suppression of oscillations in the relaxation dynamics, and in the reduction of the relaxation time.
@@
where the bar indicates normalization, $\langle v_{x}^{2}\rangle_{0}$ is the (stationary) velocity variance, with the angular brackets representing an ensemble averaged, and $\omega^{2}\equiv\Omega^{2}-\gamma^{2}/4$.
Fitting the ACF to Eq. 1 provides system-specific values of the effective inverse timescales: the damping rate, $\gamma$, and the harmonic frequency, $\Omega$.
These inverse timescales are displayed in Fig. 3F as a function of $N_{b}$, showing a transition between underdamped ($\Omega<\gamma/2$) towards critically-damped ($\Omega\approx\gamma/2$) dynamics as $N_{b}$ increases. A reduction in the relaxation time, $2\gamma^{-1}$, corresponds to an increase in collision frequency, evaluated by image analysis as the frequency of tracer-bbot collisions, $\tau_{c}^{-1}$ (black line).
At high densities ($N_{b}=15$ and $17$), fast sequential collisions require a higher temporal resolution to capture the tracer’s inertial dynamics. In these cases, we perform and use additional recordings with $\Delta t=1/60$ s to extract the effective inverse timescales (solid lines). At this sampling rate the fit parameters are consistent with the critical regime, whereas $\Delta t=1/30$ s results in overdamped behavior (dashed lines). These measurements of velocity fluctuations and their corresponding system-specific inverse timescales are further discussed in the context of the kinetic temperature in Sec. 3.2.
@@ III.1 FDR temperature
where $R_{x}$ and $R_{v}$ are the mean response functions of the position and velocity observables, respectively. The correlation functions in the (long-time) unperturbed NESS are given by a position ACF, $C_{xx}=\langle x(t)x(0)\rangle_{0}$, and a position-velocity cross-correlation function, $C_{vx}=\langle v_{x}(t)x(0)\rangle_{0}$. In this context, an effective temperature $T_{\text{eff}}$ is defined as a proportionality constant.
Notably, a static FDR can be recovered in Eq. 2 at $t=0$, with $C_{xx}(0)\rightarrow\langle x^{2}\rangle_{0}$ and $R_{x}(0)\rightarrow\Delta x_{\epsilon}$, where $\langle x^{2}\rangle_{0}$ is the position variance in the unperturbed NESS, and $\Delta x_{\epsilon}=\langle x\rangle_{\epsilon}-\langle x\rangle_{0}$ is the mean displacement between the average position in the two steady states, perturbed (sub-index $\epsilon$) and unperturbed (sub-index 0). Considering that $F_{0}=\kappa\Delta x_{\epsilon}$, the effective temperature $T_{\text{eff}}$ coincides with the definition of potential energy equipartition,
@@
Thus, $T_{\text{eff}}=T_{\text{pot}}$ is always fulfilled when the system obeys the dynamic FDR Eq. 2, even in the presence of non-Gaussian steady state distributions.
Figure 4: FDR test of the full-response to a step-perturbation: Results for mean values are obtained from pure ensemble averages over 375 two-minute step-perturbation sequences (at $t_{0}=60$ s), under weak external airflow applied by a fan at an operating voltage $10$ V; correlation functions are computed in the steady state by a combined time and ensemble average. A. Example of a typical perturbation sequence $x(t)$, where the fan is turned on for one minute and abruptly turned off for the subsequent minute. B. Time-dependent mean velocity $\langle v_{x}(t)\rangle$ and mean position $\langle x(t)\rangle$. The shown interval captures the transient response following perturbation arrest (red line) and the unperturbed steady state (a), from which correlation functions are computed. The perturbed steady state is displaced from the trap center by $\Delta x_{\epsilon}=2.2$ cm (b). C. FDR analysis using Eq. 2 (lower panel) and Eq. 3 (upper panel). These results were obtained with $N_{b}=10$. D. Parametric plot of position FDRs for $N_{b}=\{2,3,4,6,10,12,15,17\}$ realizations, where the effective temperature $T_{\text{eff}}$ determines the linear slope. In equilibrium-like behavior, all FDR data collapse onto a single line with slope $T_{\text{eff}}\sim\kappa\langle x^{2}\rangle_{0}$. Clear deviations from this equality are observed for the lowest densities ($N_{b}=2$ and 3, triangles), indicating FDR violations.
@@
Fig. 4C presents the corresponding FDR tests for both position and velocity observables. The system’s full response to the abrupt force arrest at time $t_{1}=t-t_{0}$ is evaluated as $R_{x}(t_{1})=\langle x(t_{1})\rangle-\langle x\rangle_{0}$ and $R_{v}(t_{1})=\langle v_{x}(t_{1})\rangle-\langle v_{x}\rangle_{0}$, and compared with the unperturbed correlation functions, $C_{xx}$ and $C_{xv}$. Within measurement error, these results validate the FDRs given in Eqs. 2 and 3, with an effective temperature determined by Eq. 4. Thereby, in this system the tracer’s potential temperature satisfies a dynamic FDR in both position and velocity observables. We note that the signal-to-noise ratio in $R_{v}$ diminishes with increasing $N_{b}$, rendering the velocity response experimentally inaccessible for larger values ($N_{b}>10$). The individual FDR tests of the experimental setups are further provided in the SI (see figures 3-4 therein).
Fig. 4D presents a parametric plot of the position FDR, showing normalized fluctuation and response quantities across different $N_{b}$ configurations. Specifically, the solid linear line represents validation of Eq. 2 with $T_{\text{eff}}=T_{\text{pot}}$. Violations of Eq. 2 are mainly observed with the lowest $N_{b}=2$ and $3$, indicating a breakdown of the effective temperature description (as detailed in Ref. [7]). In contrast, with larger $N_{b}$, the tracer’s potential temperature $T_{\text{pot}}$ satisfies a dynamic FDR and is therefore consistent with $T_{\text{eff}}$.
Figure 5: The modified kinetic temperature: Data are presented for configurations with $N_{b}=\{3,4,6,10,12,15,17\}$ bbots, averaged over 375 one-minute tracer trajectories recorded in unperturbed conditions. A. Comparison of the standard kinetic temperature $T_{\text{kin}}$ (solid gray line), calculated using the tracer mass $m\approx 1$ g (Eq. 5), the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (solid red line), evaluated using Eq. 6, and the potential temperature $T_{\text{pot}}$ (dashed blue line). Error bars indicate standard deviations. Dot-dashed red lines show deviations in $\tilde{T}_{\text{kin}}$ arising from increasing the measurement time interval $\Delta t$. B. Separate contributions to $\tilde{T}_{\text{kin}}$: (upper panel) the velocity variance $\sigma_{v}^{2}=\langle v_{x}^{2}\rangle_{0}$, and (lower panel) the effective mass $\tilde{m}$, evaluated from stationary ACF dynamics using Eqs. 1 and 6. Gray dot-dashed lines indicate deviations due to increasing $\Delta t$ in both panels.